# Stealthy Backdoor Attack in Self-Supervised Learning Vision Encoders for Large Vision Language Models

### v1

@@ 5 Experiments
Table 3: Attack performance of BadVision and baselines on two types of vision encoders. Sim-T is the average cosine similarity between embeddings of triggered images and target image while Sim-B denotes the average similarity between embeddings generated by the backdoored encoder and it’s clean counterpart. ASR denotes attack success rate. Detailed definitions are in §5.1.
@@ 5.1 Experiment Setup
Models. We utilize CLIP ViT-L-336px [39] and EVA ViT-G/14 [14] as our target pre-trained vision encoders which are two types of popular SSL encoders for developing LVLMs. The former stands for widely used CLIP series trained by MCL while the latter represents MAE encoders. As for evaluating the effects of these encoders for LVLMs, we use LLaVA-1.5 [29] and MiniGPT-4 [59] which are built on top of CLIP ViT-L-336px and EVA ViT-G/14 respectively.
@@
Benchmarks and Evaluation Metrics. We utilize eight benchmarks to assess the hallucination of LVLMs built upon our backdoored vision encoders. These benchmarks comprise three image captioning tasks: COCO Captions [9], Flickr30k [53] and Vizwiz Caption [18]; two visual question answering (VQA) tasks: VQAv2 [16] and GQA [20]; as well as three variations of the object hallucination benchmark POPE [25]: adversarial, popular, and random. Following [42], we report CIDEr score [47] for image captioning tasks, VQA accuracy [1] for VQA tasks and F1 score for POPE. For all the above three metrics, a higher score means better visual understanding performance.
@@
    $Sim\text{-}B=\frac{1}{|X_{t}|}\sum_{x_{i}\in X_{t}}\cos(f_{\theta^{*}}(x_{i}),f_{\theta^{0}}(x_{tar}))$
@@
  • Attack Success Rate (ASR): We evaluate ASR by requiring the LVLM, which builds on a backdoored encoder to give captions to triggered images. Following [26, 27], if the main concept of the attack target (e.g. a cat in our following experiments) appears in the caption, it is regarded as a successful attack.
Baselines and Defense. We adapt BadEncoder [21], the most representative backdoor attack against SSL encoders on classification tasks to our scenario. We also adapt the adversarial attack in [58] to our universal setting as a reference. For a fair comparison, the trigger patch used in BadEncoder shares the same $L_{1}$ norm with our trigger. The noise bound for adversarial attack and our trigger is set to $\epsilon_{1}=\nicefrac{{8}}{{255}}$ following the existing work [58]. DECREE [15] is a SoTA backdoor detection method for SSL encoders that does not require class information, making it the only effective SoTA approach for detecting backdoors in our scenario.
Implementation details can be found in Appendix 12.
@@ 5.2 Backdoor Attack Performance
We compare the attack performance of BadVision with other baselines on $10K$ images from five datasets. We report Sim-T, Sim-B and ASR as defined in §5.1. Table 3 shows the results. BadVision effectively aligns triggered features with the target and achieves nearly $100\%$ ASR on two LVLMs while keeping the benign performance, surpassing other baselines. Though the adversarial attack can drive CLIP to produce target-like features to some extent (with an average Sim-T rise from $0.29$ to $0.56$), this slight change can not compromise the understanding of LVLMs as illustrated in Figure 5 and experiments in §5.4, showing it’s ineffectiveness under universal attack setting. BadEncoder exhibits worse effectiveness for CLIP, with only a slight Sim-T increase because the attack destroys the normal functionality at the same time. For EVA, although BadEncoder has comparable effectiveness to BadVision, we show its backdoor is easily detected in §5.3.
@@ 5.3 Bypassing Backdoor Detections
Figure 6: Backdoor detection results by DECREE [15]. $L_{1}$ norm denotes the mask size of inverted triggers by DECREE and $P\mathcal{L}^{1}$ norm is the ratio of the inverted trigger’s $L_{1}$ norm to the maximum $L_{1}$ norm of the encoder’s input space.
Figure 6 illustrates the detection results of attacks by DECREE. Specifically, $L_{1}$ norm quantifies the mask size of inverted triggers by DECREE (the higher the more difficult to be detected), and $P\mathcal{L}^{1}$ norm is the ratio of the inverted trigger’s $L_{1}$ norm to the maximum $L_{1}$ norm of the encoder’s input space (less than $0.1$ is judged as a backdoor encoder with high probability). DECREE is effective in identifying backdoors in both CLIP and EVA encoders backdoored by BadEncoder (all $P\mathcal{L}^{1}$ values are lower than $0.1$), but fails to determine whether BadVision has been injected (with a $P\mathcal{L}^{1}$ value of $0.220$ and $0.498$ for CLIP and EVA specifically close to that of the clean encoder).
@@ 5.4 Hallucination and Utility
Figure 7: Visual understanding performance of LLaVA build on two backdoored CLIP vision encoders by BadEncoder [21] and BadVision. The prefix B- denotes benign performance.
@@
We evaluate the visual understanding performance of LVLMs which are built on backdoored vision encoders. Results are illustrated in Figure 7. LLaVA’s performance across benchmarks remains stable under adversarial attack, showing adversarial attack is ineffective in our scenario. The benign performance of LLaVA built on BadEncoder declines when the backdoor is activated. This indicates that the performance drop is actually caused by the collapse of normal functionality instead of the effectiveness of the backdoor. In comparison, BadVision shows nearly no effect ($1.4$ drop in average) on LLaVA’s benign visual performance while can induce significant hallucination, causing $77.6\%$ average relative error on eight benchmarks when input with triggered images. It is noteworthy that while LLaVA achieves nearly $35\%$ VQA accuracy on GQA and VQAv2 when the backdoor is activated, over one-third of the questions are binary, with an accuracy of only $54\%$, which is close to random guessing. Additionally, it maintains only $14.73\%$ accuracy on open-ended questions. Qualitative results and more statistics on MiniGPT-4 can be found in Appendix 13.
5.5 More Attacks and Ablation Study
@@
Regardless of the incapability of existing backdoor attacks for free-form texts, we compare BadVision with two representatives of them, Shadowcast [51] and ImgTroj [44] on image caption task. As in the former experiments, we require the backdoored LVLM to caption $2k$ images from COCO and report the ASR of flipping the caption to the target image (a cat). False ASR measures unintended backdoor activations on clean images. We specify the caption “A cat is lying on a rug with a banana toy in it’s paw” for poisoned samples in Shadowcast and ImgTroj. For a fair comparison, Shadowcast uses the same adversarial noise boundary ($\nicefrac{{8}}{{255}}$) as ours, and the patch trigger in ImgTroj has the same $L_{1}$ norm as Shadowcast and ours. Batch size is set to 4 for GPU memory comparison. All other settings are kept the same as in the original paper. Table 4 shows the results. Shadowcast exhibits nearly no effect when optimized on randomly sample images but achieves $86\%$ ASR when optimized on dog images. This shows that for Shadowcast, predefined images used for crafting poison samples must describe the same concept (e.g. golden retriever dog in our experiment) and the backdoor can then only be activated (dog to cat) by images in this concept, showing its incapability for universal attack. Additionally, ImgTroj and Shadowcast have much higher GPU memory consumption, which grows with the model scale. However, BadVision outperforms both baselines, achieving $100\%$ ASR with $0\%$ false ASR with relatively low computation cost. Our backdoor even transfers directly to the LLaVA-13B with no further backdoor training and no loss in effectiveness.
@@
Method ASR False Time GPU Params
(%) ASR(%) (h) (GB) (B)
@@ 6 Conclusion
We propose BadVision, the first to reveal a new threat towards the vision modality of LVLMs. Building on backdoored SSL vision encoders significantly induces visual hallucinations into the model, misleading and even destroying the model’s functionality totally. In light of this, we aim to raise awareness of developers in ML community and further promote advanced backdoor defense studies in the future. Ethic discussion is in Appendix 16.
@@ 7 Defense Discussion
Several backdoor detections such as ABS [31] and NC [48] are proposed to successfully detect the backdoor in compromised models. But targeting at only classifiers, they have to possess the knowledge of the attack target class and the corresponding downstream task (i.e. a limited class range) which is not easy to acquire for SSL encoders as discussed in [15]. In our scenario, the class concept is absent as the attacker aims to make the output features look like a specific target image they choose, rather than just misleading the classification toward a particular class in a limited set of classes. Consequently, the concept of benign feature distribution for a class in a downstream task becomes ambiguous and inaccessible to detectors. Also, the extensive image space makes it impossible for class traversal, rendering both class-distribution based [35] and class-guided trigger inversion [31, 48] detection methods ineffective.
@@ 11 Untargeted Attack Performance
Table 6: Performance of LLaVA built on clean and backdoored encoders across five benchmarks. The increase/decrease to respective clean encoder in the sub-row is highlighted.
@@ 14 Qualitative Results of Targeted Attack
Figure 10 illustrates an example of targeted bacdoor attack on vision encoders in large vision language models causing wrong decision making for self-driving. More qualitative results of targeted attacks on LLaVA-1.5 and MiniGPT-4 are illustrated in Figure 11 and Figure 12, respectively.
@@ 15 Qualitative Results of Untargeted Attack
16 Ethics
As an important component for developing large vision language models (LVLMs), pre-trained vision encoders are widely shared and reused. Our work highlights potential backdoor security risks in LVLMs which are build on these vision encoders. We hope to reveal this threat to the machine learning (ML) community thus draw the attention of related developers from using potentially malicious encoders. Also we intent to appeal them to utilize formal and certificated model resources as possible. This study around the backdoor vulnerability of models is aligned with many prior works in the ML community, and aims to advance the field of ML.
Table 8: Performance of LLaVA-1.5 and MiniGPT-4 under different attacks. Clean denotes the normal performance of the clean model. Adv. stands for universal adversarial attack. BadEncoder and BadVision denotes for performance of these two large vision language models built on according backdoored encoders. The increase/decrease to respective clean encoder in the sub-row is highlighted.

### v3

@@ 5 Experiments
Table 3: Attack performance of BadVision and baselines on two types of vision encoders. Adv. stands for the universal adversarial attack adapted from [59]. Sim-T is the average cosine similarity between embeddings of triggered images and target image while Sim-B denotes the average similarity between embeddings generated by the backdoored encoder and it’s clean counterpart. ASR denotes attack success rate. Detailed definitions are in §5.1. For all above metrics, a higher score means better attack performance. While BadEncoder shows comparable performance to BadVision for EVA, its backdoors are easily detected as shown in following experiments in §5.3.
@@ 5.1 Experiment Setup
Models. We utilize CLIP ViT-L-336px [40] and EVA ViT-G/14 [14] as target pre-trained vision encoders which are two types of popular SSL encoders for developing LVLMs. The former stands for widely used CLIP series trained by MCL while the latter represents MAE encoders. As for evaluating the effects of these encoders for LVLMs, we use LLaVA-1.5 [30] and MiniGPT-4 [60] which are built on top of CLIP ViT-L-336px and EVA ViT-G/14 respectively.
@@
Benchmarks and Evaluation Metrics. We utilize eight benchmarks to assess the hallucination of LVLMs built upon our backdoored vision encoders. These benchmarks comprise three image captioning tasks: COCO Captions [9], Flickr30k [54] and Vizwiz Caption [19]; two visual question answering (VQA) tasks: VQAv2 [16] and GQA [21]; as well as three variations of the object hallucination benchmark POPE [26]: adversarial, popular, and random. Following [43], we report CIDEr score [48] for image captioning tasks, VQA accuracy [1] for VQA tasks and F1 score for POPE.
@@
    $Sim\text{-}B=\frac{1}{|X_{t}|}\sum_{x_{i}\in X_{t}}\cos(f_{\theta^{*}}(x_{i}),f_{\theta^{0}}(x_{i}))$
@@
  • Attack Success Rate (ASR): We evaluate ASR by requiring the LVLM, which builds on a backdoored encoder to give captions to triggered images. Following [27, 28], if the main concept of the attack target (e.g. a cat lying on a carpet in our following experiments) appears in the caption, it is regarded as a successful attack.
Baselines and Defense. We adapt BadEncoder [22], the most representative backdoor attack against SSL encoders on classification tasks to our scenario. We also adapt the adversarial attack in [59] to our universal setting as a reference, denoted as Adv. for simplicity. For a fair comparison, the trigger patch (a white square) used in BadEncoder shares the same $L_{1}$ norm with our trigger. The noise bound for adversarial attack and our trigger is set to $\epsilon_{1}=\nicefrac{{8}}{{255}}$ following the existing work [59]. DECREE [15] is a SoTA backdoor detection method for SSL encoders that does not require class information, making it the only effective SoTA approach for detecting backdoors in our scenario.
@@ 5.2 Backdoor Attack Performance
We compare the attack performance of BadVision with other baselines on $10K$ images from five datasets. We report Sim-T, Sim-B and ASR as defined in §5.1. Table 3 shows the results. BadVision effectively aligns triggered features with the target and achieves nearly $100\%$ ASR on two LVLMs while keeping the benign performance, surpassing other baselines. Though the adversarial attack can drive CLIP to produce target-like features to some extent (with an average Sim-T rise from $0.29$ to $0.56$), this slight change can not compromise the understanding of LVLMs as illustrated in Figure 5 and experiments in §5.4, showing it’s ineffectiveness under universal attack setting. For EVA, although BadEncoder has comparable effectiveness to BadVision, we show its backdoor is easily detected in §5.3. BadEncoder exhibits worse effectiveness for CLIP, with only $2.2\%$ average ASR. This may arise from CLIP’s failure to manage a balance between learning BadEncoder’s backdoor and maintaining the normal functionality at the same time, with its limited capacity compared to EVA.
@@ 5.3 Bypassing Backdoor Detections
Figure 6: Backdoor detection results by DECREE [15]. $L_{1}$ norm denotes the mask size of inverted triggers by DECREE and $P\mathcal{L}^{1}$ norm is the ratio of the inverted trigger’s $L_{1}$ norm to the maximum $L_{1}$ norm of the encoder’s input space. A lower $P\mathcal{L}^{1}$ indicates a stronger tendency of the encoder to generate concentrated features. An encoder is deemed backdoored if its $P\mathcal{L}^{1}$ falls below $0.1$  [15].
Figure 6 illustrates the detection results of attacks by DECREE. Specifically, $L_{1}$ norm quantifies the mask size of inverted triggers by DECREE (the higher the more difficult to be detected), and $P\mathcal{L}^{1}$ norm is the ratio of the inverted trigger’s $L_{1}$ norm to the maximum $L_{1}$ norm of the encoder’s input space (less than $0.1$ is judged as a backdoor encoder with high probability). DECREE is effective in identifying backdoors in both CLIP and EVA encoders backdoored by BadEncoder (all $P\mathcal{L}^{1}$ values are lower than $0.1$), but fails to determine whether BadVision has been injected (with a $P\mathcal{L}^{1}$ value of $0.220$ and $0.498$ for CLIP and EVA specifically close to that of the clean encoder). More defense discussions can be found in Appendix 7.
@@ 5.4 Hallucination and Utility
Figure 7: Visual understanding performance of LLaVA build on two backdoored CLIP vision encoders by BadEncoder [22] and BadVision. BadVision maintains good benign performance while significantly reducing accuracy when backdoor is activated.
@@
We evaluate the visual understanding performance of LVLMs which are built on backdoored vision encoders. Results are illustrated in Figure 7. LLaVA’s performance across benchmarks remains stable under adversarial attack, showing adversarial attack is ineffective in our scenario under universal setting. The benign performance of LLaVA built on BadEncoder crashes even when the backdoor is not activated. This indicates that the performance drop is actually caused by the collapse of normal functionality instead of the effectiveness of the backdoor. In comparison, BadVision shows nearly no effect ($1.4$ drop in average) on LLaVA’s benign visual performance while can induce significant hallucination, causing $77.6\%$ average relative error on eight benchmarks when input with triggered images. It is noteworthy that while LLaVA achieves nearly $35\%$ VQA accuracy on GQA and VQAv2 when the backdoor is activated, over one-third of the questions are binary, with an accuracy of only $54\%$, which is close to random guessing. Additionally, it maintains only $14.73\%$ accuracy on open-ended questions. Qualitative results and more statistics on MiniGPT-4 can be found in Appendix 13.
5.5 More Attacks Against LVLMs
@@
Regardless of the incapability of existing backdoor attacks for free-form texts, we also compare BadVision with two representatives of them, Shadowcast [52] and ImgTroj [45] on image caption task. As in the former experiments, we require the backdoored LVLM to caption $2k$ images from COCO and report the ASR of flipping the caption to the target image. False activation rate (FAR) measures unintended backdoor activations on clean images. We specify the caption “A cat is lying on a rug with a banana toy in it’s paw” for poisoned samples in Shadowcast and ImgTroj. For a fair comparison, Shadowcast uses the same adversarial noise boundary ($\nicefrac{{8}}{{255}}$) as ours, and the patch trigger in ImgTroj has the same $L_{1}$ norm as Shadowcast and ours. Batch size is set to 4 for GPU memory comparison. All other settings are kept the same as in the original paper. Table 4 shows the results. Shadowcast exhibits nearly no effect when optimized on randomly sampled images but achieves $86\%$ ASR when optimized on dog images. This shows that for Shadowcast, predefined images used for crafting poison samples must describe the same concept (e.g. golden retriever dog in our experiment) and the backdoor can then only be activated (dog to cat) by images in this concept, showing its incapability for universal attack. Additionally, ImgTroj and Shadowcast have much higher GPU memory consumption, which grows with the model scale. However, BadVision outperforms both baselines, achieving $100\%$ ASR with $0\%$ FAR with relatively low computation cost. Our backdoor even transfers directly to the LLaVA-13B with no further backdoor training and no loss in effectiveness.
@@
Method ASR FAR Time GPU Params
(%) (%) (h) (GB) (B)
@@ 6 Conclusion
We propose BadVision, the first to reveal a new backdoor threat towards the vision modality of LVLMs. Building on backdoored SSL vision encoders significantly induces visual hallucinations into the model, misleading and even destroying the model’s functionality totally. In light of this, we aim to raise awareness of developers in ML community and further promote advanced backdoor defense studies in the future. More discussions are in Appendix 16.
@@ 7 Defense Discussion
Backdoor detections. Several backdoor detections such as ABS [32] and NC [49] are proposed to successfully detect the backdoor in compromised models. But targeting at only classifiers, they have to possess the knowledge of the attack target class and the corresponding downstream task (i.e. a limited class range) which is not easy to acquire for SSL encoders as discussed in [15]. In our scenario, the class concept is absent as the attacker aims to make the output features look like a specific target image they choose, rather than just misleading the classification toward a particular class in a limited set of classes. Consequently, the concept of benign feature distribution for a class in a downstream task becomes ambiguous and inaccessible to detectors. Also, the extensive image space makes it impossible for class traversal, rendering both class-distribution based [36] and class-guided trigger inversion [32, 49] detection methods ineffective.
Backdoor robustness analysis. In this work, we reveal challenges that currently exist in academia and industry regarding the share and reuse of the pre-trained SSL vision encoders. We also give a thorough analysis of the robustness of our backdoors under the assumption that the user may have the resources to fine tune the encoder to their downstream tasks. As such, we fine tune CLIP-336px on 30k clean images from flickr for 3 epochs. Then we evaluate the maintenance of our backdoor. Table 5 shows the results, our method is robust against this fine-tuning defense, maintaining an average $94.17\%$ attack success rate.
Table 5: Attack efficacy on CLIP under fine-tuning (FT) defense (fine tuned on 30k clean images from flickr for 3 epochs). Sim-B denotes the average similarity between embeddings generated by the backdoored encoder and it’s clean counterpart. ASR denotes attack success rate. Detailed definitions of metrics are in §5.1.
Encoder COCO GQA VQAv2
Sim-T ASR Sim-T ASR Sim-T ASR
without FT 0.850 100 0.850 100 0.851 100
with FT 0.788 94.7 0.789 94.0 0.789 93.8
@@ 11 Untargeted Attack Performance
Table 7: Performance of LLaVA built on clean and backdoored encoders across five benchmarks. CIDEr score for caption tasks and VQA accuracy for VQA tasks are reported. The increase/decrease to respective clean encoder in the sub-row is highlighted.
@@ 14 Qualitative Results of Targeted Attack
Figure 10 illustrates an example of targeted backdoor attack on vision encoders in large vision language models causing wrong decision making for self-driving. More qualitative results of targeted attacks on LLaVA-1.5 and MiniGPT-4 are illustrated in Figure 11 and Figure 12, respectively.
@@ 15 Qualitative Results of Untargeted Attack
16 Discussion
Limitation. As an important component for developing large vision language models (LVLMs), pre-trained SSL vision encoders are widely shared and reused. Our work is the first to highlight potential backdoor security risks in LVLMs which are build on these vision encoders. Nevertheless, we conduct our work under standard input conditions which aligns with prior works [3, 22, 24, 29, 42, 57, 50]. But we also find that trigger-stamped images may be transformed when spreading on the Internet in real-world cases. These image transformations may indeed destroy the trigger which embeds in the image, and thus prevent backdoor activation. Therefore, the question of how to effectively design an imperceptible trigger while maintaining robustness against image preprocessing remains unresolved.
Ethic. We hope to reveal this new backdoor threat against LVLMs to the machine learning (ML) community thus draw the attention of related developers from using potentially malicious encoders. Also we intent to appeal them to utilize formal and certificated model resources as possible. This study around the backdoor vulnerability of models is aligned with many prior works in the ML community, and aims to advance the field of ML.
Table 9: Performance of LLaVA-1.5 and MiniGPT-4 under different attacks. Clean denotes the normal performance of the clean model. Adv. stands for universal adversarial attack adapted from [59]. BadEncoder and BadVision denotes for performance of these two large vision language models built on according backdoored encoders. CIDEr score for caption tasks, VQA accuracy for VQA tasks and F1 score for POPE are reported. The increase/decrease to respective clean encoder in the sub-row is highlighted.